OpenStudy (anonymous):

5.Solve the equation for the given unknown. Show full solution for full credit. 6x + 9/5x+3=-6

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY!! :D

OpenStudy (anonymous):

thanks please help

OpenStudy (akashdeepdeb):

As the Code-Of-Conduct here puts it well - I cannot give you the solution. But I can give you the bits and peices so that you can form your own solution!

OpenStudy (anonymous):

WELCOME to open study !

OpenStudy (anonymous):

i just need the answer :/

OpenStudy (anonymous):

LOoK!: you can have + with x s like them .

OpenStudy (akashdeepdeb):

\[\frac{ 6x+9 }{ 5x+3 } = -6\] NOW CROSS MULTIPLY IT \[6x+9 = -6(5x+3)\] NOW OPEN THE BRACKETS \[6x + 9 = -30x - 18\] Now TAKE THE x's TO ONE SIDE AND SOLVE THE SIMULTANEOUS EQUATION!! :D Understood? :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wats the answer

OpenStudy (anonymous):

Solve the equation 6x+9 / 5x+3 = -6 1. multiple 5x+3, both side 2. so now 6x+9 = -6(5x+3) 3. 6x+9 = -30x-18 4. 36x=-27 5.x= -27/36 = -3/4

OpenStudy (akashdeepdeb):

See @brauza This is your question not mine you have to find out the answer yourself Go through this http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

ty iaru

OpenStudy (anonymous):

6x+9/5 = 15/5x You get it first .

OpenStudy (anonymous):

Now we have : 15/5x+3=_ 6

OpenStudy (anonymous):

so its -3/4

OpenStudy (anonymous):

What ?

OpenStudy (anonymous):

the answer

OpenStudy (anonymous):

And what do you want to know ?!:)

OpenStudy (anonymous):

if the answer is really -3/4

OpenStudy (akashdeepdeb):

Did you understand how to get it @brauza ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

k thx

OpenStudy (anonymous):

I was saying to you .. Your answer is true ...!:)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can u help with another?

OpenStudy (anonymous):

another ?!

OpenStudy (anonymous):

another question

OpenStudy (anonymous):

Yeah ! Look:

OpenStudy (anonymous):

look?

OpenStudy (anonymous):

3x+4=-5 x=?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Is it your means ?

OpenStudy (anonymous):

what

OpenStudy (anonymous):

The Question you want ...

OpenStudy (anonymous):



OpenStudy (anonymous):

@;shkrina : I dont want to you answer !

OpenStudy (anonymous):

oh. solve for x. x^2yz = 5k/x^5p^2

OpenStudy (anonymous):

Aha! Wait ... !

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Do you want X?!

OpenStudy (anonymous):

it just says solve for x

OpenStudy (anonymous):

What slove ?!!!

OpenStudy (anonymous):

What do you mean ?

OpenStudy (anonymous):

thats all it says :(

OpenStudy (anonymous):

Can you ask your problem again ?!

OpenStudy (anonymous):

solve for x. x^2yz = 5k/x^5p^2

OpenStudy (anonymous):

I say the thinks . I wish it s your answer :

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

We can have : x^7.p^2.y.z=5k

OpenStudy (anonymous):

@E.ali give her the ans ... wahy r u making it so long ..

OpenStudy (anonymous):